OpenStudy (anonymous):

A system of equations is shown below. 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (wolf1728):

Basically, you want another system of equations that will have the same solution as the original. Why not just multiply both equations by 2?

OpenStudy (wolf1728):

8x + 5y = 9 3x + 2y = 4 x=-2 y = 5 Multiplying both equations by 2 16x + 10y = 18 6x +4y =8 x = -2 y = 5

OpenStudy (anonymous):

im being yelled at

OpenStudy (secret-ninja):

^random. O.o @Brie624 please tell me you weren't murdered...

OpenStudy (anonymous):

i am very much alive lol @Secret-Ninja